### v1

# The WINTER Observatory: A One-Degree InGaAs Survey Camera to study the Transient Infrared Sky

Abstract
The Wide-field Infrared Transient Explorer (WINTER) is a near-infrared time-domain survey instrument installed on a dedicated 1-meter robotic telescope at Palomar Observatory. The project takes advantage of the recent technology advances in time-domain astronomy, robotic telescopes, large-format sensors, and rapid data reduction and alert software for timely follow up of events. Since June of 2023, WINTER robotically surveys the sky each night to a median depth of mag, balancing a variety of science programs including searching for kilonovae from gravitational-wave alerts, blind surveys to study galactic and extragalactic transients and variables, and building up reference images of the near-infrared sky. The project also serves as a technology demonstration for new large-format Indium Gallium Arsenide (InGaAs) sensors for wide-field science in the near infrared without cryogenically cooled optics or detectors. WINTER’s custom camera combines six InGaAs sensors with a novel tiled fly’s-eye optical design to cover a 1 degree-squared field of view with 1 arcsecond pixels in the -, -, and shortened--band filters (0.9 - 1.7m). This paper presents the design, performance, and early on-sky science of the WINTER observatory.
I Introduction
Despite the recent rise in optical time-domain surveys, such as ATLAS (atlas), DeCAM (DECAM), Pan-STARRS (pan), ZTF (ztf2019), and GOTO (goto) and the upcoming Rubin Telescope’s LSST survey (Rubin), the near-infrared time-domain sky remains relatively unexplored. High sky backgrounds and the cost of the state-of-the-art HgCdTe (Mercury Cadmium Telluride) sensors prevent similar growth in near-infrared time-domain astronomy. Legacy and contemporary near-infrared surveys—for example, TMSS (tmss), 2MASS (2mass), UKIRT (UKIRT), VISTA (VISTA), WISE/NEOWISE (wise), and Palomar Gattini IR (De:2019xhw)—often trade off between coverage and depth, resulting in either relatively shallow surveys (for example, Palomar Gattini IR, mag; De:2019xhw) or are restricted to small areas on sky (such as the UKIRT Infrared Deep Sky Survey (UKIDSS) which covered 7,500 square degrees (UKIRT) or the VISTA Variables in the Via Lactea (VVV) survey covering 520 square degrees (VISTA)). Consequently, in near-infrared wavebands, there is a gap in deep, all-sky reference images and characterization of near-infrared transients which would benefit the time-domain community and lay groundwork for the upcoming Nancy Grace Roman Space Telescope.
The Wide-Field Infrared Transient Explorer (WINTER) is a new near-infrared time-domain survey instrument on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie:2020; Frostig:2020). WINTER covers over 1 square degree with a 90% fill factor in the (1.0 m), (1.2 m) and shortened -bands (1.6 m). The instrument was commissioned in June of 2023 and is operating robotically each night, surveying around 400 degrees per week to an average depth of magnitudes.
WINTER is designed around three primary goals: (1) rapid robotic follow-up of gravitational-wave alerts to search for kilonovae, (2) near-infrared time-domain surveys, and (3) general observing support for the WINTER community (Frostig:2020). The near-infrared colors uniquely study the r-process materials produced in kilonovae, allowing two weeks or more to search for electromagnetic signatures from the gravitational-wave events (Frostig:2022). Beyond kilonovae, near-infrared observations are well suited for dust-obscured, cool, and intrinsically red transients, variables, and static sky science. Between following up gravitational-wave alerts, WINTER scans the accessible northern sky in a series of Galactic and extragalactic surveys studying a range of science including transiting exoplanets around low-mass stars, tidal disruption events, supernovae, stellar mergers, and young star outbursts (See Section II for a detailed introduction to WINTER science cases). Over a ten-year nominal lifespan, WINTER will also build up a deep, co-added image of the near-infrared sky, preparing for future infrared surveys and studying galactic structure, high-redshift quasi-stellar objects, and brown dwarfs.
In addition to WINTER’s science drivers, the instrument also demonstrates the feasibility of Indium Gallium Arsenide (InGaAs) hybridized focal plane arrays as a cheaper alternative to HgCdTe sensors for near-infrared astronomy. Unlike HgCdTe, InGaAs sensors do not require cryogenic cooling to achieve background limited performance from the ground, further reducing the cost and complexity of the project. A prototype camera demonstrated this approach with InGaAs sensors and achieved sky-background limited photometry while running around -40∘C with a thermoelectric cooler (TEC) and secondary liquid cooling (Simcoe:2019aps; Sullivan:2013). WINTER also features a novel “fly’s-eye optical” design to fill the 1 square degree field of view with six non-buttable sensors with only minimal gaps, and has custom readout electronics controlled by field programmable gate arrays (FPGAs), enabling fine control over the highly-programmable sensors.
WINTER is leading a new class of InGaAs-based astronomical surveys, which currently includes the upcoming DREAMS telescope in the Southern Hemisphere (DREAMS). WINTER, DREAMS, and the newly operating HgCdTe-based PRime-focus Infrared Microlensing Experiment (PRIME; PRIME_performance; PRIME_camera; PRIME_optics), are at the head of a new effort to expand our knowledge of the near-infrared time-domain sky.
This paper describes the WINTER instrument and early performance since its installation in June 2023. A more detailed treatment on WINTER’s science cases can be found in Section II. Section III describes the InGaAs sensors and their performance, along with WINTER’s custom readout electronics, firmware, and software for controlling the sensors. The instrument design is described in Section IV, including the camera’s novel fly’s-eye optical design, the mechanical design, optomechanics, and cooling of the instrument. Section V describes the observatory and 1-meter robotic telescope and Section VI outlines the robotic control, scheduling, and data reduction software. Finally, the paper discusses on-sky performance including a brief description of current surveys and preliminary results from select early science investigations (Section VII), and a conclusion (Section VIII).
II Science Objectives
WINTER’s design directly flows down from its science goals: including the need for rapid robotic follow up with low overheads, near-infrared coverage, and a wide field of view (Frostig:2020). The science cases highlighted in this section introduce the motivation behind building the instrument, but are not a comprehensive list.
II.1 r-process elemental synthesis
Near-infrared observations are uniquely suited for studying kilonovae, thermal transients powered by rapid neutron-capture (r-process) nucleosynthesis in the neutron-rich ejecta from binary neutron star (BNS) or neutron star black hole (NSBH) mergers. The landmark BNS event, GW170817, detected in both gravitational waves and across the electromagnetic spectrum, demonstrated how heavy element synthesis in ejecta abundant in r-process materials blankets optical wavelengths, shifting the afterglow into the NIR (TheLIGOScientific:2017qsa; GBM:2017lvd; Kasen:2017sxr). Unlike optical emissions, which are model-predicted to be short-lived (under one week) and highly viewing-angle dependent, NIR emission is both long-lived (over one week) and isotropic, making it a promising kilonova signature for BNS and some NSBH mergers (Kasen:2013xka; Barnes:2016umi). Moreover, NIR surveys could detect up to 8–10 times more kilonovae than optical surveys (Zhu:2020ffa).
WINTER’s design directly supports kilonova discovery through several key features: (1) sensitivity to the Lanthanide-rich kilonova emission peak in the near-infrared; (2) a wide field of view; (3) fast robotic follow-up with low overhead resulting in the ability to quickly tile gravitational-wave uncertainty contours. The need for efficient BNS merger follow-up, which yielded no detections during the third gravitational-wave observing run (O3), informed WINTER’s project requirements, including the optical design and ensquared energy criteria, the noise budget and instrument cooling, and optical performance which minimize false positives in wide-field searches (Frostig:2020).
Frostig:2022 outlines an end-to-end simulation of WINTER follow-up observing of BNS mergers during the O4. The study predicts NIR kilonovae are longer lived and red kilonovae are detected 1.5 times further in the infrared than in the optical. The study provides a detailed simulation of WINTER’s projected ability to discover new kilonovae, estimating up to five new discoveries in O4. However, the sensitivities of both WINTER and the international network of gravitational wave detectors were lower than expected throughout O4: no BNS mergers were detected by LIGO, and despite several international follow-up campaigns of BBH and BH-NS mergers, there were no confirmed optical counterparts of GW events (e.g., Pillas:2025, including WINTER data).
II.2 Peering through the dust
Heavily obscured regions can conceal infrared transients and variable sources. Programs like the SPitzer InfraRed Intensive Transients Survey (SPIRITs), which surveyed 200 nearby galaxies, identified 2457 variables and 78 transients over four years, with 60 transients too red for optical detection, suggesting that optical surveys may miss up to 40% of core-collapse supernovae (CCSNe) in dusty environments (Kasliwal_SPIRITs_2018; Jencson_2019). SPIRITs also uncovered a new class of intermediate-luminosity transients, bridging the gap between novae and supernovae (Kasliwal_SPIRITs_2017). Infrared time-domain observations also reveal variability in young stellar objects (YSOs) across diverse environments; for instance, Spitzer found that 50% of YSOs in regions like the Orion Nebula and Large Magellanic Cloud are variable, enabling studies of jets, accretion disks, and circumstellar matter (yso_mag; yso_orion; yso_var). Infrared surveys are equally important for mapping obscured Galactic structures, as demonstrated by the VVV survey, which uncovered previously unknown variable stars and star-forming regions, while UKIDSS expanded our knowledge of faint stellar populations in the Galactic plane (VISTA; UKIRT).
II.3 Studying cool objects
Infrared observations are particularly effective for studying cool objects such as brown dwarfs, planets, red giants, and asymptotic giant branch (AGB) stars, as these objects have lower temperatures (below 300 K for brown dwarfs up to 5,000 K for red giants) that peak in infrared wavelengths (Faherty; Marocco). Infrared time-domain surveys have significantly advanced our understanding of red giants and AGB stars, revealing pulsation periods and episodic mass-loss events observable through dust production and circumstellar changes (Liljegren). Long-Period Variables (LPVs), including Miras and semi-regular AGB stars, display complex variability patterns, which WISE and VVV surveys have explored in depth (Groenewegen; Nikzat). For brown dwarfs, Spitzer’s time-resolved spectroscopy and photometry have shown rotational modulations related to cloud patterns and temperature variations, enriching atmospheric models (Metchev; Artigau). Infrared studies also illuminate cool object binaries, where interactions with compact companions, such as in symbiotic systems, produce infrared variability through accretion processes, jet formation, and dust in outflows (Miko). Furthermore, exoplanet studies around M-dwarfs, the most common stars in the Milky Way, benefit from infrared observations, as these cool, low-mass stars offer favorable conditions for detecting transits in red optical to infrared bands (Dressing; Charbonneau). For example, the WINTER prototype InGaAs camera has observed exoplanet transits at Las Campanas Observatory, showcasing the potential of infrared tools for exoplanet discovery (Simcoe:2019aps).
II.4 Static sky science
Throughout the course of WINTER’s nominal ten-year lifespan, the project aims to build up a deep reference image of the near-infrared northern sky. In addition to facilitating transient detection, utilizing this reference image to find new sources in nightly WINTER images, the reference image will help build a near-infrared map of L and T dwarfs, quasars, and galactic structure while also providing a reference image for the broader community. The static reference image and a long-baseline time-domain survey will also provide valuable preparation for the upcoming time-domain surveys from the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope.
III Sensors, readout electronics, firmware, and software
III.1 InGaAs sensors
The six InGaAs sensors central to WINTER are custom AP1020 focal plane arrays (FPAs) developed in collaboration with Teledyne-FLIR and Sensor Creations (now Attollo Engineering Inc./Safran Defense & Space Inc.) for our program (Figure 2). With a high-definition format of 1920 by 1080 pixels and a 15m pitch, the sensors are the largest format InGaAs sensors ever manufactured for commercial use. InGaAs exhibits dark current 2–3 orders of magnitude lower than 1.7 m cutoff HgCdTe at equivalent temperature (Seshadri_2007_MCT_vs_ingaas_JPL_2007) down to C, or in the 170–300 K range. For ground-based imaging applications like WINTER, adequate dark current to reach background limited performance can thus be reached without cryogenic cooling, and the 1.7 m cutoff reduces sensitivity to thermal backgrounds compared to detectors like 2.5 m cutoff HgCdTe which require cryogenically cooled optics.
The hybridized complementary metal oxide semiconductor (CMOS) focal plane arrays consist of photosensitive lattice-matched InGaAs PIN (p-type/insulator/n-type) diodes hybridized to a readout integrated circuit (ROIC) with a capacitive transimpedance amplifier (CTIA) pixel architecture. With a lower-energy bandgap than silicon, InGaAs is sensitive in the near/shortwave-infrared (NIR/SWIR) between 0.9 m and its 1.7 m room-temperature direct bandgap cutoff wavelength. The highly-programmable ROIC allows for multiple imaging modes, field-programmable gain, and application-specific sensor tuning. In contrast to the source follower per detector (SF) readout common in astronomy detectors like the Teledyne Hawaii family, the CTIA architecture supports video rate imaging at the price of increased read noise (ROIC_pixel_architectures_guellec_2014). An optimized CTIA readout can also deliver a higher degree of linearity than SF. To help mitigate this increased read noise (see Section III.4.4), the AP1020 ROIC features options for on-chip correlated double sampling and multiple non-destructive reads which substantially reduce the effective read noise. The sensor read out with eight interleaved data output channels, allowing for full frame, continuous sampling at a frame rate of up to 30 Hz. The focal plane array is bonded to a two-stage thermoelectric cooler (TEC) in hermetically-sealed vacuum housing (see further details in Section III.3).
III.2 Readout
Custom electronics, firmware, and software are required to control the InGaAs sensors and to accommodate WINTER’s size, heat, and readout speed requirements. The sensors are individually managed by dedicated electronics controlled by FPGAs, sending data over fiber to a computer located in a nearby shed to minimize heat buildup in the dome. This configuration allows for low-noise and low-latency data acquisition within a relatively compact, freely rotating instrument. The custom readout scheme also leverages the fast readout of the sensors. Reading each pixel continuously throughout an exposure–sampling multiple times per second–enables noise reduction through common methods such as Fowler pairing and sample-up-the-ramp. Furthermore, the fast readout software works in tandem with robotic observing to enable near-real-time transient detection and alerts. Further discussion on the WINTER readout design can be found in Malonis:2020 and Frostig:2022SPIE.
III.2.1 Electronics
WINTER’s focal plane electronics are split into five boards, assembled into a focal plane module for each detector (Figure 2):
1.
The sensor board plugs directly into the InGaAs sensor and serves as a motherboard, connecting the other boards to the sensor with minimal signal path lengths.
2.
The analog board converts the image data to digital signals with eight high-speed, 16-bit analog to digital converters (ADCs).
3.
The FPGA module is an Opal Kelly XEM7310 board with a Xilinx Artix 7 FPGA, controlling the other boards and the InGaAs sensors and providing image clocking, housekeeping, and power control.
4.
The digital board interfaces between the FPGA module and the other electronics.
5.
The power board distributes power to the sensor and electronics and hosts a thermoelectric cooler (TEC) drive circuit and a housekeeping system for closed-loop temperature control.
III.2.2 Firmware and software
The FPGA firmware runs the sensors, TEC, and focal plane readout electronics locally and interfaces with the WINTER control software. The firmware is written in Verilog, fully modeled in Vivado Simulator, and provides three main functions: electronic control, sensor control, and interfacing with the software. To operate the sensors, the firmware provides clocking, imaging signals, and writes the highly programmable ROIC settings, enabling imaging in various modes at a range of frame rates. Moreover, the firmware controls the power sequencing, telemetry, and housekeeping, and data collection of the sensors and readout electronics along with a safety watchdog overseeing system temperatures and connectivity to the software. Finally, the firmware communicates with the software through a USB 3.0 connection using the Opal Kelly Host Interface, providing an interface to enact commands from the software and to collect and stream sensor data.
The six FPGA modules stream data over USB 3.0, converted to OM4 multimode fiber, to a readout computer in a neighboring shed. Each sensor has its own Icron USB 3.0 Spectra 3022, allowing full bandwidth data link while extending the range of the USB allowing the readout computers to be located outside of the telescope dome and providing electrical isolation of the USB link to avoid ground loops in the readout electronics. An independent multi-threaded, python-based software daemon independently controls each sensor in parallel for rapid data readout and temperature monitoring, and a top level daemon coordinates the image acquisition on each sensor and collates the data into a single multi-extension FITS file.
III.2.3 Read-out modes
The WINTER sensors feature two implemented read-out modes: integrate while read (IWR) and non-destructive read (NDR). IWR is a simpler read out mode where a single image is sampled at the end of the exposure. In this case, the data are read out while the subsequent image is exposing to reduce overheads. This simpler mode was deployed during commissioning of the WINTER sensors and is the current operating mode, since exposures are not read-noise limited.
Alternatively, NDR allows for a reduction in read noise (see Section III.4.4) and is an available mode for observing with WINTER, although it is not currently used in the survey. Below we describe the read-out scheme for NDR.
For short exposures, data are streamed in at seven frames per second, utilizing every available frame, and for longer exposures, the FPGA streams in every other frame, pausing between frames to control temperature deviations.
At the end of an exposure, the continuously sampled frames are immediately pre-processed into single-frame, six-sensor FITS format images which are then downlinked from the observatory for further data reduction (see Section VI.3). To pre-process the WINTER data, the many readout frames are collapsed into one image with one of two techniques: (1) correlated-double sampling (CDS), in which the final frame is subtracted from the first frame, or (2) sample-up-the-ramp (SUTR), in which a linear fit is used to get a slope of the signal in each pixel. SUTR reduces read noise (see Section III.4.4) and allows for recovering data affected by saturation, cosmic rays, and streaming errors, but also adds significant data volumes and overhead time.
To facilitate transient discovery, WINTER requires rapid data pre-processing and reduction to minimize data storage volume and enable real-time decisions on candidate events and observing strategies. The use of graphics processing units (GPUs) is well-suited to efficiently process high-definition imaging data. WINTER’s readout computer has two commercial NVIDIA GeForce RTX 3080 Ti GPUs, each one processing three sensors’ images in parallel. For ease of use, the GPU code uses cupy, an open-source GPU-implemented python library modeled after numpy. Implementing the nonlinearity correction and NDR linear fit on GPUs can lead to a factor of 10 speed up from a CPU-only implementation for a five minute exposure with six sensors.
III.3 Thermal control
WINTER’s InGaAs sensors use a vacuum-sealed package with a two-stage thermoelectric cooler (TEC) at the core of their cooling system. Each sensor has its own mechanical and thermal module which includes a kinematic mounting system for holding the detector allowing manual focus adjustments during assembly, and provides the required cooling to the detectors TEC assembly. An oxygen free high conductivity (OFHC) copper sled, sandwiched between the back of the sensor and the readout electronics, draws heat away from the sensor with two sintered copper/water heat pipes leading to a liquid heat exchanger, placing any liquid outside of the main instrument box and away from the optics and electronics (Frostig:2022SPIE). The heatpipes soldered with solder paste into a OFHC copper flange which is bolted to a small liquid heat exchanger with an indium foil interface material to maximize conductive coupling. An OptiTemp OTC 1.0A closed-loop 70%/30% distilled water/propylene glycol mix chiller located in a shed next to the telescope dome provides parallel cooling to each of the six focal plane modules through a manifold located on the telescope rotator. The chiller is set to 10∘C which maintains the warm side of the TEC to 12∘C through the 15 m of hose between the chiller and the instrument. The sensor readout computer controls the TEC setpoint with a proportional-integral-derivative (PID) loop. The instrument itself is sealed with a series of gaskets to keep out dust, pollen, and other environmental contaminants common at Palomar Observatory. To avoid condensation within the camera enclosure during humid days, a dry air purge system consisting of a 1 horsepower Powerex oil-less compressor and a nano purification systems dryer delivers air with a -40∘C dewpoint. This system delivers 28 liters per minute of air which is piped through the camera with a series of 1/4 inch hoses with small holes drilled in them, which provides sufficient airflow through the box to keep critical electronics like the FPGA modules from overheating.
WINTER is designed for sky-background-limited photometry, with TEC-based cooling to reducing the dark current levels below the sky background. The smaller prototype instrument employed a single-stage TEC and demonstrated background-limited performance between -C to -C, with dark current halving for every 7∘C of cooling (Simcoe:2019aps). However, for the manufactured WINTER sensors, a combination of TEC under-scope and insufficient process control in sensor packaging resulted in decreased cooling capacity and varied final operating temperatures across the six sensors (between -10∘C and -30∘C). Section III.4.3 provides a detailed discussion of the sensor cooling.
III.4 Performance
The WINTER project initiated the first manufacturing run of these sensors, incorporating a new ROIC design. The vendor manufactured the sensors on a best-effort basis and enough sensors were produced to fill in WINTER’s focal plane; however, the sensors did not meet several key performance specifications, as discussed in this section. Overall, this deviation from expected performance results in a factor of 10 reduction in instrument efficiency (2 magnitudes decrease in depths) from the instrument design. In total, fourteen sensors were received, with varying levels of functionality. Six of the highest-performing sensors were integrated into WINTER, with five non-responsive units allocated for thermal testing and three kept as spares.
| Metric | Predicted | Measured | Units |
| Gain | 1 (spec.) 1.6 (design) | 2.5 | e-/DN |
| Quantum efficiency | 0.80 | 0.05–0.10 | |
| Dark current | 113 | 120–370 | e-/s/pix |
| Read noise | 45 | 50–70 | e- |
III.4.1 Conversion gain
Characterizing the WINTER sensors begins with measuring the conversion gain that links photoelectron signal to the corresponding digital output. Based on the ROIC design, the gain is estimated to be g=1.6 e-/DN (see Frostig:thesis for the full calculation). Typically, this value is experimentally verified using a Photon Transfer Curve (PTC) measurement, which assumes the presence of a shot-noise limited regime where noise increases with signal level N as . Multiple PTC approaches explored in Frostig:thesis, reveal nonlinear responses (see Section III.4.5) and differing gain estimates for the NDR and IWR read out modes. Each method produces a gain estimate within a error of the expected g=1.6 e-/DN. Furthermore, the gains vary by sensor, between the eight read out channels within each sensor, and across the pixels within those read out channels.
For simplicity, we present one measurement of the gain here and use it throughout the paper, but note the significant spread in gain by pixel, requiring careful flat-fielding. Figure 4 shows the range of slopes from a PTC for one readout channel of a WINTER sensor, taken with a set of laboratory flats on an integrating sphere. The variance is taken for a stack of ten dark-subtracted frames at each exposure time, with eight rounds of sigma clipping to reject pixels outside of the three standard deviations of the mean. A line is fit to the slope of the variance vs mean signal for each pixel. This analysis predicts a mean gain of g=2.5 e-/DN but shows a spread of responses across pixels within this read out, with some pixels consistent with the designed gain of g=1.6 e-/DN. Given this uncertainty, we present all analyses throughout this study in both measured counts (DN) and converted to electrons (e-).
III.4.2 Quantum efficiency
We measured a quantum efficiency (QE) of less than 10% for the WINTER sensors, significantly lower than the designed 80%, indicating substantial losses in photon-to-electron conversion efficiency. This measurement was validated through three independent tests: (1) laboratory characterization of the full system using a controlled light source, (2) direct laboratory measurement of the sensors’ photocurrent response, and (3) on-sky calibration with reference stars. Laboratory measurements were made using a 20 inch integrating sphere illuminated with a laser driven light source from Energetiq and a series of optical and near infrared narrow bandpass filters. The absolute optical flux was determined using a silicon photodiode with a NIST-traceable responsivity calibration, allowing the quantum efficiency to be measured between 500 and 1050 nm. The results revealed nonlinear behavior (see Section III.4.5) and a peak response of approximately 5% QE, suggesting a manufacturing issue leading to deviations from the expected efficiency (Figure 6) Furthermore, in standard QE and photon transfer curve (PTC) measurements, conversion gain and QE are inherently degenerate, necessitating an independent test to pinpoint the root cause of the discrepancy.
To isolate the sensor photodiode array’s intrinsic response to light from any effects introduced by the readout integrated circuit (ROIC) or the custom readout chain, we conducted a secondary set of QE tests. These tests employed an external source measurement unit to directly supply the sensor bias voltage, allowing precise voltage control and direct photocurrent measurement on the sensor while bypassing the readout circuitry via programmable logic on the ROIC. This approach enabled a direct measurement of photocurrent response, confirming the measurement of a peak QE of (Figure 6). We define the quantum efficiency as the ratio of the number of electrons generated per incident photon. Given an input optical power (in nW) and a measured photocurrent (in µA), the QE can be expressed as:
where is the measured photocurrent in amperes (A), is the incident optical power in watts (W), is Planck’s constant (J·s), is the speed of light (m/s), is the elementary charge (C), and is the wavelength of the incident light (m).
To further evaluate the sensor’s response, we consider the relationship between the measured photocurrent and the expected photocurrent from perfect charge transfer (Figure 6). The expected photocurrent under ideal conditions, , is compared to the actual measured photocurrent . The effective quantum efficiency can be determined from the linear regression slope in the relation:
where represents the QE, and accounts for any systematic offset in the measurement. Figure 6 shows the results of this testing for three spare WINTER sensors, one of which (SCI022) was on-sky for one year. The three sensors show QEs between 6% and 10%, significantly lower than the specified design of % in all bands.
Finally, on-sky flux observations of standard stars independently measure WINTER’s full instrument and telescope throughput. Accounting for as-measured reflectivities of the mirrors and lenses, the results confirm an average sensor QE of 10 (see Frostig:2024 for a detailed calculation). This analysis includes sensor serial number SCI022, which was previously in the Starboard B position in the full instrument, which has a laboratory measured QE of 9.6% (Figure 6).
The decreased QE when directly probed on the sensors most likely arises from an issue in the manufacturing or quality control during hybridization, photodiode array production, or processing. The test procedures confirms that poor QE does not result from deficiencies in the ROIC or amplification/read-out chain. Since the InGaAs growth and photodiode array design made use of mature processes at Teledyne-FLIR, we conclude that this production lot of a new device experienced issues in preparation and bonding of InGaAs to the Si multiplexer, a hypothesis that could be verified with destructive testing and microscopy of a sensor from the lot.
III.4.3 Dark current
Figure 7 displays the dark current in the six WINTER sensors by temperature in the final integrated camera on the telescope. The sensors reach a minimum temperature between T C and T C, unable to achieve the designed cooling of T C. The temperature discrepancy arises from two key factors. First, the TEC selection process did not include realistic modeling of the sensor packaging, resulting in the use of an undersized two-stage TEC (Laird MS2-107-10-10-12-12-11-W8). Second, the final design placed an excessive heat load on the TEC through parasitic radiative coupling and thermal conduction through wire bonds from the pins. Temperature variations between sensors are influenced by differences in TEC bonding, vacuum seal integrity, and the temperature of each sensor board, which impacts the TEC’s hot-side temperature. The dark current decreases approximately linearly with the logarithm of the temperature down to about C, beyond which it plateaus, possibly due to thermal infrared emission from the detector housing.
Figure 7 also highlights the sensor-to-sensor variation in dark current among WINTER’s six on-sky sensors. Part of the apparent spread reflects the C uncertainty in the on-chip temperature calibration, but intrinsic differences are also present. Port C—consistently the coldest device—exhibits roughly a factor of six lower dark current than Port B at their respective operating temperatures. These relative background levels are visually apparent in the focal-plane image shown in Figure 3.
III.4.4 Read noise
One key advantage of the WINTER sensors over other InGaAs detectors is the ability to read out the sensors non-destructively throughout an exposure. This allows for read-noise reduction through algorithms like sample-up-the-ramp (NDR), recovery of saturated data, and real-time elimination of cosmic rays (Robberto_2014). The non-destructive read mode is not currently enabled in nightly WINTER operations. This section explores the read noise measured with and without non-destructive read.
To evaluate the read noise in the WINTER sensors we analyze the reference rows of each sensor. These reference rows, positioned outside the active 1920 × 1080 photosensitive region, are insensitive to incident radiation. As a result, any signal variations in these pixels arise predominantly from read noise without incorporating the effects of pixel response nonlinearity (see Section III.4.5) or photon shot noise. Figure 8 illustrates an analysis of the signal standard deviation for each reference pixel over 384 frames in a non-destructive read ramp. The median read noise for a single readout, equivalent to the current on-sky IWR mode, is approximately 24.5 DN RMS (equivalent to 61.3 e-, assuming a gain of 2.5 e-/DN), which is close to the initial project requirement of 45 e- RMS, depending on the exact gain per pixel.
By continuously reading out the sensors with non-destructive sampling, multiple exposures of the same pixel can be averaged before a reset, reducing the random electronic noise in accordance with the theoretical scaling. Figure 8 demonstrates how this technique lowers the effective read noise as frames are binned in increasing groups from two to 64. The standard deviation of the readout signal along a best-fit line decreases, with the lowest read noise reaching approximately 3 DN (7.5 e-) when averaging in groups of 164 frames.
Despite achieving read noise levels within project specifications, even with a single read, a low-frequency read-noise component introduces a time-dependent striped pattern in the images. This structured noise is effectively removed in the data reduction pipeline using Fourier filtering (see Section VI.3).
III.4.5 Linearity
WINTER’s sensors exhibit stronger nonlinearity than originally expected, introducing additional challenges for photometry of faint sources (Figure 9). The most pronounced deviations occur at low to moderate signal levels, with measurable curvature beginning near 10,000 counts and continuing well into the mid-well regime. Even in the region of highest linearity, roughly 30,000–45,000 counts, the response is not fully linear.
To correct for these effects, we derive a per-pixel piecewise polynomial calibration. Each pixel is modeled with an 11th-order polynomial for the low-count region and a 5th-order polynomial for the high-count region, joined at a fixed crossover point of 25,000 counts with a 10,000-count overlap region. The fitting uses uniformly resampled exposure-time sequences of dome flats. During science processing, raw counts are passed through the appropriate segment of the piecewise model to yield an approximately linearized response, as seen in the dashed lines of Figure 9. The nonlinearity correction results in a median nonlinearity % across the WINTER sensors. The effect of these corrections on the science data is explored in V. Karambelkar and R. D. Stein et al., in prep.
To make a bad-pixel mask, a line is fit to the central 33–85% of each pixel’s dynamic range—the portion of the well that behaves most linearly—and residuals relative to this baseline are evaluated across the full well. Pixels whose residual standard deviations exceed 500 ADU, including unresponsive, unstable, and hot pixels, are flagged and added to the bad-pixel mask.
The nonlinearity correction coefficients, bad pixel masks, generation and implementation code are all released publicly through the winternlc package available on git.111https://github.com/winter-telescope/winternlc
IV Instrument Design
IV.1 Optical fly’s-eye design
WINTER’s optical design covers a 1 square degree field of view with 90% fill factor despite the sensors’ vacuum-sealed housing preventing close tiling (Figure 2). The novel fly’s-eye optical design splits the image between six identical optical channels to direct each section of the image onto its respective sensor with minimal inter-chip gaps between channels (Figure 10). To enable close packing of the optics, each lens is truncated symmetrically about one plane, matching the aspect ratio of the HD-format sensors. After passing through a waveband filter, the telescope’s focal plane is divided into six channels by a field-lens array made of six bonded lenses. A right-angle fold mirror then divides the image into two halves, placing three sensors on either end of the instrument to keep the center of gravity close to the telescope rotator (Hinrichsen:2020).
Furthermore, WINTER’s optical design prioritizes a 1″ pixel scale for the 15 m sensor pitch to match the median seeing at Palomar Observatory and other time-domain surveys such as the Palomar Transient Factory (PTF) and the Zwicky Transient Facility (ZTF). This necessitates a 2:1 demagnification of the F/6 telescope beam to an F/3 beam at the sensor. Each optical channel consists of ten spherical lenses. Lenses 1-5 collimate the telescope F/6 beam onto a pupil between lenses 5 and 6, and then lenses 6-10 reimage the pupil at an F/3 beam onto its respective InGaAs sensor.
All lens groups are bonded into groups of three tiled lens elements, with the exception of the field lens array (lens 1), which has six bonded lenses, and the lens 6-7 doublet, which retains independent adjustment to act as a compensator lens for fine focus optimization. Tolerancing studies simulating alignment procedures for the WINTER optical system for using Zemax’s Monte Carlo simulations found that maintaining strict fabrication tolerances for the lenses allows for loosening of mechanical alignment tolerances (Lourie:2020; Frostig:2020). This approach simplifies the optical alignment procedure, allowing for vertically tiled lens elements to be aligned together with the lens 6-7 bonded doublet allowing for independent adjustment for each optical channel. All WINTER lenses were purchased from Optimax Systems, who co-developed the precision truncation and alignment of the bonded, tiled lens elements, and all glasses were purchased from Ohara Corporation from the same melt batch to reduce variation between optical channels. To verify infrared performance, an outside company, M3, measured the index of refraction of a sample of each glass from 0.7 – 1.7 m. The infrared properties of these glasses had not yet been reported in the literature and closely matched predicted performance.
IV.2 Filters
WINTER observes with , , (shortened H) filters based on the Mauna Kea Observatory filter set with a long-wave cutoff in H-band at 1.7 m to match the InGaAs bandgap cutoff. The filters, manufactured by Asahi Spectra on 10 mm fused silica, achieve 99 in-band throughput while blocking out-of-band light to 0.01 (Figure 11). The filters are placed 50 mm in front of the instrument field lens array. Placement within the telecentric, F/6 converging beam of the telescope ensures spectral uniformity across the field of view, simplifying photometric calibration, with a modest reduction in the slope of the waveband edges due to the angle of incidence variation within the F/6 beam.
IV.3 Filter tray
WINTER’s aluminum filter tray, which holds three filters (, , and ) and a mirror for dark frames, is mounted to the front of the instrument, perpendicular to the optical axis (Figure 10). A stepper motor with a pinion is coupled to a gear rack and actuates each filter change. The position of each of the filters is marked by a detent, which in conjunction with a spring-loaded arm, ensures the stability and stationarity of the filter tray when in use. Additionally, a linear encoder attached to the tray monitors the filter tray position to correct for any slippage that may occur during operation. Upon startup each night, the filter tray is homed to the dark mirror and each subsequent filter change between adjacent filters takes 15 seconds.
IV.4 Optomechanics
The WINTER optomechanical design employs adhesive bonding for the lenses to reduce mass, design complexity, and stress on the optics. 3M Scotch-Weld 2216 A/B Gray epoxy was selected for all adhesive bonds based on extensive thermal and humidity lifetime tests simulating WINTER’s ten-year nominal lifetime (Hinrichsen:2020). The vertically-tiled lens elements are mounted in MIC-6 aluminum frames with flexure mounts made of Titanium 6Al-4V, providing single-axis lateral adjustment of the lens group (Figure 12). A custom 5-axis adjustable mount, made from Titanium 6Al-4V and Aluminum 7075-T6, holds the compensator lenses (6-7 doublet) with a stacked stage design where each stage adjusts one axis. The adjustment stages are accessible outside of the main instrument enclosure to provide flexibility for future focus modifications.
The right-angle fold mirror is the heaviest optic in the instrument and the original adhesive bond design failed during initial alignment. Instead, the fold mirror has a six-point kinematic mount that fully constrains all degrees of freedom. The mount is loaded against six hardpoints with spring plungers, four on the front optical surfaces, one on the back face, and one on the top. To reduce the contact stresses, the spring plungers and hard points are all spherical contact surfaces with a machined Delrin cap to spread the load. All optomechanics are rated for a 6g shipping load, although the fold mirror was removed from the optical assembly during shipping.
To mitigate stray light, the optomechanical design also includes a focal plane mask and baffles around the tiled lens groups (Figure 12). The focal plane mask is placed in front of the field lens to prevent cross-talk between optical channels and reflections off of the bonded edges of the tiled lenses, which was identified as a major source of stray light in the non-sequential ray-tracing simulations outlined in (Frostig:2020). The focal plane mask is milled from MIC-6 cast aluminum and painted with Krylon Ultra-Flat Black, with the openings matching the footprints of the sensors on the telescope focal plane. Similar aluminum baffles painted with Aeroglaze Z306 flat black are mounted to the tiled lens group frames, with openings oversized relative to the image footprint. Prototype focal plane masks made with selective laser sintering (SLS) 3D printers were produced as well. The 3D printed versions captured the complex geometry of the baffle but did not achieve the knife-edge sharpness of the machined masks. While these were not optically characterized, with further optimization this approach may be viable for future efforts.
IV.5 Alignment, integration, and test
WINTER’s custom truncated lenses require special alignment, integration, and test (AIT) procedures. In this section, we provide a concise overview of the AIT procedure, while a comprehensive analysis of the instrument AIT process and outcomes will appear in an upcoming publication. First, the as-manufactured lens thicknesses were measured with a TriOpics OptiSurf interferometer, prompting a re-optimization of the lens spacings based on the tolerancing procedures outlined in Lourie:2020; Frostig:2020. Next, the fold mirror was installed in the center of the instrument with a Hexagon coordinate measuring machines (CMM) arm used to verify the placement and tilt of the optic to a 25 m accuracy. Small mirrors were attached to the sensor mounts on either end of the instrument and aligned with a Zygo Verifire wavefront interferometer to pre-align the tip, tilt, and focus position of the sensors before delivery from the vendor. The tiled lens group were then aligned with the CMM arm, with the aluminum frames resting against reference pins and shimmed to compensate for tip and tilt along the optical axis. The individually-mounted compensator lenses were then aligned with their five-axis mounts and verified with the CMM arm. For the final alignment, a fiber-fed pinhole mask manufactured by FiberTech Optica provided an array of five F/3 sources per sensor, allowing for fine optical three-axis positioning of the sensors and verifying the required instrument image quality. Due to time constraints and the initial alignment meeting design requirements, a final optical alignment with compensator lenses was not conducted.
IV.6 Camera housing/mechanics
A precision optical bench serves as the reference plane for aligning and mounting all optomechanical components, including the sensor mounting frames (Figure 13). The optical bench is made of a cast aluminum plate with flatness and parallelism tolerances of 0.025 mm and placement tolerance for mounting holes and alignment pins of 0.18 mm. It is painted with Aeroglaze Z306 flat black paint between non-critical surfaces to reduce stray light. Black-painted panels, resting on an aluminum frame and sealed with rubber gaskets to prevent dust contamination, form the remaining five sides of the enclosure. A focal plane mask located in front of the field lens array mitigates stray light due to cross-talk between optical channels (Figure 12). Similar baffles around the tiled lens arrays act as stray-light reduction and a safety retaining system in case of adhesive bond failure (Figure 12). To further prevent stray light near the image planes, Acktar Advanced Coatings’ Metal Velvet Black Foil was applied to the sensor mounting frames and parts of the filter tray. A custom telescope adapter made from carbon fiber reinforced plastic manufactured by CarbonVision mounts the instrument to the telescope rotator.
V Observatory
WINTER operates at Palomar Observatory in Southern California, sharing a site with the 200-inch Hale Telescope, the Zwicky Transient Facility (ZTF) (ztf2019) and Palomar Gattini IR (De:2019xhw). The site is at an altitude of 1700 meters and is run by Caltech Optical Observatories. WINTER takes advantage of existing infrastructure at the site to run robotic telescopes, including data downlink, weather monitoring, and facilities operations. The weather and observing conditions are continuously monitored by telescope operators and support astronomers at the 200-inch Hale Telescope, who enable and lockout observing throughout the night accordingly. This continuous staffing support also allows for prompt hands-on response to operational anomalies and equipment failures that may occur over years of operation.
The instrument operates in a preexisting five-meter dome with a new, adjacent climate-controlled shed housing the WINTER control computers and electronics, power supplies, and chiller (Figure 15). The dome was retrofit for robotic operations, including a robotic control and telemetry system similar to those developed for ZTF and PGIR, a new weather station, motor drives, and a custom telescope pier. The dome’s robotic control system allows for extensive autonomous safety protocols to protect both workers and equipment, as outlined in Frostig:2020.
V.1 Telescope
WINTER employs a 1-meter commercial off-the-shelf (COTS) corrected Dall-Kirkham (CDK) telescope made by PlaneWave Instruments with modifications for the project (Figure 15). The PW1000 1-Meter Observatory System features an instrument rotator mount to correct for sky rotation and fast direct-drive altitude and azimuth motors with a maximum velocity of 15 degrees per second with a settle time between slews of one second. The telescope optics include a 1-meter primary mirror, a spherical secondary mirror, a flat tertiary mirror with an integrated rotator to select between the two Nasmyth ports, and a set of three corrector lenses at each of the ports. All of the telescope optics are made of fused silica to increase thermal stability. Customizations for the WINTER project include developing a focusing system at the secondary mirror instead of the default focus mechanism at the instrument rotator mount, and applying near-infrared-optimized anti-reflection coatings on the corrector lenses for one Nasmyth port. The opposing port retains optical-wavelength-optimized coatings and houses a companion instrument.
V.2 SUMMER and SPRING
The telescope and the companion instrument, SUMMER (Studying the Universe with Multi-Messenger and Exoplanet Research), were installed in June of 2021, ahead of WINTER’s completion in 2023, enabling development of the robotic control code, scheduler, and data reduction pipeline (see sections VI.1, VI.2, VI.3) and performing optical follow up of science targets (e.g., Christos:2023). The instrument consists of a Raptor Photonics CCD with e2v 42-40 architecture, which includes a 5-stage TEC and liquid cooling loop, a 65 mm optical shutter from Vincent Associates, and a Finger Lakes Instrumentation seven-position filter wheel which includes u, g, r, and i SDSS passband filters from Asahi Spectra. SUMMER images cover a 0.29-degree-squared field of view.
Since commissioning the telescope, the second port on the 1-meter telescope also served as a testbed for new sensors. There have been multiple observing runs with optical Complementary Metal-Oxide-Semiconductor (CMOS) sensors, including the Hamamatsu Photonics qCMOS camera and the large-format Teledyne COSMOS camera (see results in Layden:2025). Currently, it is home to the SPRING (Super PalomaR INGaas) camera, a Princeton Infrared Technologies 1280SciCam InGaAs camera with 1280 x 1024 12.5m pixels used for deep infrared observations of targets that fit within the smaller 8.8 x 7.0 arcminute field of view.
VI Software
VI.1 Robotic control software
The WINTER telescope’s operations are managed by the WINTER Supervisory Program (WSP), a highly parallelized Python 3.9-based control system leveraging PyQt5222https://riverbankcomputing.com/software/pyqt/ for interfacing with the Qt application framework333https://www.qt.io/development/qt-framework, enabling sophisticated control and scheduling capabilities (wsp_software_spie) 444https://github.com/winter-telescope/observatory/. The WSP consists of more than a dozen individual subsystem daemons which communicate with each other over TCP/IP sockets using the Pyro5 python library555pyro5.readthedocs.io, each of which are multi-threaded to allow concurrent commanding, communication, and housekeeping monitoring. The system is launched and managed as a Linux system daemon on the observatory’s control PC, with individual subsystem daemons run on separate PCs. WSP generates nightly schedules and manages observation tasks. This system, drawing upon methodologies from the MINERVA (MINERVA_Swift_2015), Robo-AO (robo-ao_Riddle_2012), GOTO (goto_dyer) and SuperBIT (superbit_romualdez) projects, facilitates automated decision-making for observatory operations, such as dome opening and observation initiation, while allowing manual override by operators in case of adverse conditions. The WSP orchestrates a variety of processes including scheduling, target of opportunity (ToO) event handling, command execution via a priority queue system, and comprehensive logging of telescope states and observations through real time telemetry graphing and slack notifications.
VI.2 Scheduler and ToOs
To autonomously balance multiple observing programs with a range of cadences, filters, and priorities, WINTER employs a custom scheduling software based on the ZTF scheduler (ztf_scheduler_Bellm_2019). It creates nightly schedules based on WINTER’s surveys and long-term targets, which can be interrupted by Target of Opportunity (ToO) requests at any point in the night. A flexible API, winterapi666https://github.com/winter-telescope/winterapi allows authorized users to submit ToO requests directly to the queue, download data associated with their observing programs, and enables high-level tracking by the team to ensure adequate use of observing time. This framework allows observers to manage their own programs, and enables automated follow-up triggering of transient candidates without the need for an observer in the loop. The scheduler prioritizes programs based on their observational history such that any observations interrupted by weather or ToOs are rescheduled for subsequent nights. The WINTER scheduler divides the accessible northern sky ( declination) into a fixed pointing grid of fields matching the WINTER field of view. Each potential observation is prioritized by a volumetric weighing metric and then scheduled with one of three modes: Gurobi-optimized observing, greedy observing, and queue observing. The volumetric metric maximizes the limiting volume of a requested field at a given altitude, which is based on detailed simulations of the limiting magnitude including seeing, sky brightness, and instrument noise. For nightly surveys, Gurobi-optimized observing utilizes the Gurobi linear optimizer to solve a traveling salesman problem, balancing various scientific surveys with unique cadences while minimizing overhead time from telescope slews. Alternatively, greedy observing continually selects the best field based on current volumetric weighting and is used to fill in reference sky images between scheduled science surveys and ToOs. Finally, for any ToO requests, queue observing sequentially steps through a predefined program. Figure 16 shows WINTER observations from the first two years of observing, balancing various surveys as discussed in Section VII.
For robotic follow up of gravitational-wave alerts, WINTER uses the Growth ToO Marshall (GROWTH) to ingest alerts distributed via the LIGO-Virgo GCN system and produce optimized observing schedules with the gwemopt package (Coughlin:2018lta; Coughlin:2019qkn; Ghosh:2015sxp). Frostig:2022 outlines extensive simulations of WINTER observing with the gwemopt package, leading to a tailored gravitational-wave follow up strategy based on the reported distance and localization area of each gravitational-wave event. WINTER only follows up events with at least a 10 chance of kilonova detection based on simulations with two kilonova model grids. For each qualifying event, WINTER dynamically adjusts the image exposure time to optimize the search strategy, dedicating up to seven nights searching for each event and repeatedly observing kilonova candidates until they fade.
VI.3 Data reduction pipeline
The WINTER data reduction and transient alert pipeline build upon the work for the ZTF and Palomar Gattini IR projects. The data reduction pipeline, named the Modular Image Reduction and Analysis Resource777https://github.com/winter-telescope/mirar (mirar) is a open-source, modular, Python 3.11-based pipeline (V. Karambelkar and R.D. Stein et al., in prep). mirar processes both SUMMER and WINTER images, going through a series of standard infrared data reduction steps:
1.
Quality cuts: Rejection of bad images and masking of bad sensor regions.
2.
Calibration: Dark subtraction, flat-field correction, removal of sensor pattern noise through Fourier filtering, and subtraction of the modeled sky background with the SExtractor package (Bertin1996).
3.
Astrometric calibration: Each image is astrometrically calibrated using astrometry.net (astrometry.net) and a cross-matched catalogue of sources detected in Gaia (Gaia) and 2MASS (Cutri2003). Bad images are once again rejected at this step.
4.
Stacking: Each WINTER image is split into a set of shorter exposures dithered around a central point, to enable long exposure times with the bright infrared sky background and reconstruction of the point spread function (PSF). The dithered exposures are stacked using the Swarp algorithm (swarp).
5.
Photometric calibration: Images are calibrated with Gaia, 2MASS, and Pan-STARRS.
6.
Difference imaging: Nightly images are subtracted with the ZOGY package (zogy) from UKIRT (J and Hs), Pan-STARRS (Y), or WINTER (Y, J, and Hs) reference images, where available.
7.
Candidate identification: The SExtractor package is used to create a catalog of sources from the difference images, which allows from cross-matching, filtering, and resolved source naming.
8.
Real-bogus classification: A machine-learning classifier distinguishes between image subtraction artifacts and real transient and variable sources.
Sources classified as “real” are uploaded with the Apache Avro888https://avro.apache.org/ data serialization system to IPAC at Caltech, where they are sent to the Kowalski alert broker and the Skyportal frontend (skyportal) for alert visualization and follow-up coordination (Figure 17). Candidates appear daily on the Fritz web portal, which is a specific instance of Skyportal with a Kowalski backend, for human vetting and coordination of follow-up observations.
VII On-sky Performance and Early Science
WINTER has been on sky since June of 2023, running robotically to observe a mix of long-term surveys with ToO interrupts. The supervisory program, wsp, manages scheduling, sensor startup and shutdown, focusing, science observations, and dark frame collection. All operations are logged in real time on a dedicated Slack channel, with alerts sent to relevant personnel. While a Caltech Optical Observatories member supervises each night, the system generally runs without human intervention. The data reduction pipeline processes the data in real time and candidate transient events are reviewed by members of the WINTER team.
VII.1 Photometric depth
The original design for WINTER predicted a limiting magnitude of for a 5-minute exposure under ideal observing conditions (Frostig:2020). Multiple factors in the realization of the instrument led to the limiting magnitudes of WINTER to be in a 16-minute exposure (Figure 18). The most significant change is the effective QE, reduced from the designed QE (see Section III.4.2). Additionally, increased dark current, higher gain, and a larger focused spot size compared to the original design introduce secondary performance degradations (Table 1). There are also significant variations in performance between the six sensors, caused by differences in quantum efficiency, gain, dark current at the final operating temperatures, and hybridization issues present in some sensors (e.g., the glow spots seen in Figure 3). This leads to a median depth in a 16-minute exposure of 17.6-18.6 in (8 dithers), 17.6-18.7 in (8 dithers), and 16.7-17.7 in (15 dithers) for the six sensors for MJD 60849 - 60856. The data reduction pipeline is optimized around -band imaging, with and optimization planned. The depths also vary significantly with the conditions of the sky background.
VII.2 Early science with WINTER
WINTER observations balance a mix of long-term surveys with ToO interrupts (Figure 16 shows the current history of WINTER observations). The core WINTER community surveys target regions of the sky predicted to yield the most infrared transients and variables in a blind search. This includes a 14-day cadence Galactic plane survey in - and -bands and a series of galaxy-targeted surveys. Galaxies within 25 Mpc, ULIRGS, and galaxy clusters are surveyed at a 10-day cadence in -band. For the first year of operations, very nearby galaxies, such as M31, were monitored daily in -, -, or -bands. In addition to the core surveys, WINTER builds up -band reference images of the sky, prioritizing fields above 60∘ declination where there are no UKIRT or VISTA images. The surveys are regularly interrupted by ToO requests with minutes-scale response times. The WINTER collaboration has published over seventy General Coordinates Network (GCN) circulars, including regular -band detections of gamma-ray bursts (e.g., 2025GCN.42562....1M) and detections of potential counterparts to neutrino alerts (e.g., 2024GCN.35584....1S).
WINTER sets itself apart from other infrared surveys with its time-domain infrastructure, enabling rapid alerts and prompt spectroscopic follow-up of transient sources. For example, WINTER observed a young stellar object outburst (Figure 19). The team followed up the event with the MMIRS spectrograph on the MMT (MMIRS; MMIRS_drp) and classified the event as an FU Ori type outburst following the classification scheme in Connelley:2018. This event is also observed and studied in (Gaia18dvy). WINTER also observed a highly reddened galactic classical nova (2024ATel16658....1W) (Figure 19). Follow-up spectroscopy with the FIRE spectrograph on the Magellan telescopes (FIRE) shows the presence of carbon lines, which could indicate the event is a Fe II classical nova (NIRnovae). Other science with WINTER includes infrared coverage of supernovae, which are well covered in optical surveys but not infrared surveys (Figure 20), and luminous red novae in nearby galaxies.
The collaboration has released three studies with early WINTER data. Frostig:2025 present NIR follow-up of the GW candidate S250206dm, covering 43% of the 50% localization region but finding no viable kilonova candidates, consistent with the large inferred distance of 373 Mpc. Karambelkar:2025 report the slow eruption of an early-AGB star in M31 (WNTR23bzdiq/WTP19aalzlk), which exhibited a decade-long optical/IR brightening and spectroscopic properties resembling stellar merger transients, possibly signaling the onset of common-envelope evolution. Finally, Frostig:2025b characterize two infrared FU Ori outbursts discovered by WINTER and NEOWISE, including one new confirmation of an embedded Class I protostar, demonstrating WINTER’s ability to uncover obscured protostellar eruptions inaccessible to optical surveys.
VIII Conclusions
WINTER is running nightly robotic near-IR surveys since June 2023. Through a mix of extragalactic and Galactic plane surveys, targeted follow-up, and tiled searches, the project is enabling studies of reddened and intrinsically red transients and variables. The time-domain infrastructure supports rapid follow-up of new discoveries. Recent highlights from early WINTER science include searches for kilonovae from gravitational-wave triggers, evidence of common envelope evolution in AGB stars, and characterization of outbursts and mass accretion in young stellar objects.
While WINTER’s sensors did not meet design specifications, the project demonstrates the promise and need for wide-format InGaAs sensors at an accessible cost. Extensive testing presented here found decreased quantum efficiency from the design specification (from 80% to 10%) due to an issue in hybridization or processing of this lot of sensors. The InGaAs technology has no fundamental flaws, and the read out integrated circuit, and custom read out electronics all performed near expectations. The decreased quantum efficiency, along with secondary effects from increased dark current, gain, and focus spot size, reduced the on-sky depth by 2 magnitudes. However, a broad range of science is underway at the present average depth of mag, which is significantly fainter than historical IR time domain surveys.
WINTER studies complement and prepare for other infrared surveys. Future surveys such as the Roman Space Telescope and Cryoscope (Cryoscope) will enable wider and deeper near-IR observations, extending studies of these transients to greater depths. In the meantime, WINTER is developing near-IR infrastructure, surveys, and science for the time-domain era.
IX Acknowledgments
WINTER’s construction is made possible by the National Science Foundation under MRI grant number AST-1828470 with early operations supported by AST-1828470. Significant support for WINTER also comes from the California Institute of Technology, the Caltech Optical Observatories, the Bruno Rossi Fund of the MIT Kavli Institute for Astrophysics and Space Research, the David and Lucille Packard Foundation, and the MIT Department of Physics and School of Science.
WINTER continued operation relies on the dedicated efforts of the Caltech Optical Observatories engineering and technical staff, support astronomers, and telescope operators at Palomar Observatory. Particularly, but not exclusively, Robert Sandoval, Gregord van Idsinga, Jon Derek Knapp, Drew Roderick, Isaac Wilson, Kathleen Koviac, and Paul Nied.
The team acknowledges the ongoing dedicated efforts of Stefan Lauxtermann and Per-Olov Pettersson and the detector teams at Sensor Creations/Attollo Engineering, throughout the detector production and commissioning, and Bryan Gall at Teledyne-FLIR.
Others at MIT and Caltech contributed to various aspects of the WINTER software and pipeline development, and analysis, including Benjamin Schneider, Tomas Ahumada, and Christopher Layden. Undergraduate and high school students contributed to the program with support from the Caltech SURF and MIT UROP programs: Allan Garcia-Zych, Joshua Glass, Sulekha Kishore, and Aswin Suresh.
Many others provided invaluable input to and review of various hardware and software subsystems.
D.F.’s contribution to this material is based upon work supported by the National Science Foundation under Award No. AST-2401779. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time.
This work made use of Astropy:999http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (astropy:2013; astropy:2018; astropy:2022).